OpenStudy (anonymous):

the length of a rectangle is 6 times its width. if the perimeter is 70 ft what is the area

OpenStudy (anonymous):

Let the width be x and length be y

OpenStudy (anonymous):

y = 6x . P = 2l + 2w , P = 2(6x) + 2x , P = 12x + 2x , P = 14x , 70 ft = 14x , x = 5 ft . y = 6x . y = 30 ft . width = 5 ft and length = 30 ft. Area = l * w , area = 30 * 5 , area = 150 ft squared.